# SSD: A State-based Stealthy Backdoor Attack For Navigation System in UAV Route Planning

### v1

@@ 6.4 Attack Stealthiness
We also use two UAV-specific detectors, NLC [30] and LTW [31], as baselines for comparison. These methods are highly effective at detecting GNSS attacks and they maintain both strong accuracy and response time, even when identifying targeted stealthy attacks. We follow the threshold in previous work [30]. Fig. 11 demonstrates that in the LTW test group, the detection statistic consistently remains below the threshold boundary, indicating the superior stealth characteristics of SSD attacks. Notably, NLC statistics demonstrate the following three characters (shown in Table III):
@@
These findings demonstrate significant motion-dynamic sensitivity disparities in NLC detectors. However, when attackers combine motion-state transition strategies, positional attack can realize residual normalization during linear phases and reduce detection sensitivity velocity attack counteract statistical fluctuations from nonlinear components. These two methods make SSD show greater stealthiness in hybrid motion missions.
@@
6.5 Attack Effectiveness
We start by determining the optimal combination of $Q_{k}$ and $R_{k}$ for each scenario through offline learning. With these optimal values, we proceed to validate the effectiveness of the SSD. For each combination of INS and trajectory, the UAV is tasked with following the designated path to complete a full mission, while the SSD is deployed to attack the flight. Table IV presents the changes in each metric before and after the attack. On average, ADE/FDE is increased by 425%/591%. The lateral (N)/longitude (E) deviation reaches / meters. We will analyze the factor that based on the experiment on three scenarios.
@@ 6.5 Attack Effectiveness
Normal/Attack Normal/Attack Normal/Attack
ECL EKF2 Trajectory I 20s (0.95)/(5.68) 597.89% (1.80)/(5.89) 327.22% (0.89)/(5.66) 635.96%
CD-EKF Trajectory I (1.62)/(4.64) 286.42% (1.49)/(7.54) 506.04% (1.81)/(4.92) 273.48%
ECL EKF2 Trajectory II 48s (1.31)/(4.09) 312.21 % (1.78)/(6.15) 345.51% (1.48)/(4.11) 277.70%
CD-EKF Trajectory II (1.78)/(8.53) 479.21 %
ECL EKF2 Trajectory III 71s (1.46)/(5.38) 366.44% (0.95)/(10.47) 1102.11% (1.86)/(6.49) 348.92%
CD-EKF Trajectory III (1.65)/(5.38) 326.06% (0.56)/(10.47) 1869.64
@@
(c) route
Figure 12: Attack visualization in Trajectory III
6.5.1 Different Scenarios
@@
In terms of scenarios, the SSD shows a greater increase in positioning error in the purely linear motion state (Fig. 12(a)) compared to the purely nonlinear state (Fig. 12(b)). This difference arises from the time-varying nature of the velocity vector in the nonlinear state, which leads to attenuation of the indirect positional interference caused by the velocity perturbation $G(t_{i},a^{d}_{i};\phi)$. However, experiments with Section 6.4 demonstrate that a velocity attack can still ensure the fulfillment of the stealthy precondition. When the mission involves mixed kinematic modes, the combination of velocity and positional attacks results in an additive interaction, as shown in Table IV. The baseline increase is 366.44% for ADE and 1102.11% for FDE, confirming that SSD enhances attack effectiveness synergistically. Next, we will quantitatively analyze the impact of this state on mission completion rates.
6.5.2 Impact on Mission
@@
When the UAV’s trajectory deviation exceeds the tolerance range, the mission completion rate shows a clear decline. In three different scenarios, the APDE metrics increased by $635.96\%$, $277.70\%$, and $348.92\%$, respectively, indicating that SSDs significantly compromise mission reliability. Notably, in Scenario Trajectory III, where the mission duration is 71 seconds, the FDE and APDE deviations reached $10.47m$ and $6.49m$, respectively—the largest among the three scenarios. This demonstrates that SSD lowers the mission completion rate and impedes the UAV’s ability to return accurately after completing the mission (shown in Fig. 12(c)). This effect highlights its kinematic chain impact:since the mission-critical point is typically located at the motion state transition (e.g., hover → level flight, acceleration → deceleration). The attack-induced cumulative error in position estimation propagates into the next motion state, where it is further amplified.
7 Discussion
@@
This study reveals the coupling mechanism between UAV motion dynamics and the cyber attacks effectiveness through systematic empirical analysis. Experimental data indicate that changes in motion state can significantly enhance the success rate of attacks. While the SSD approach demonstrates clear advantages, its engineering implementation faces two major challenges: Dependency on A Priori Knowledge. The effectiveness is highly contingent upon the real-time accuracy of the object detection and tracking system. The existing YOLOv5 architecture, for instance, exhibits exponential decay in the Intersection over Union (IoU) metric over time in dynamic target tracking scenarios. This results in a tracking failure probability exceeding 73% after 60 seconds of continuous locking. Energy-Concealment Trade-off Paradox. While the sustained attack mode can maintain a stealthiness threshold, it leads to a non-linear increase in energy consumption on the attacking end. This escalation doesn’t align with the requirements in real-world mission scenarios.
@@
In addition, a key assumption of the SSD is that the $Q(k)$ and $R(k)$ are preset offline. This is a common practice in current engineering applications. From a theoretical standpoint, adaptive $Q(k)$ and $R(k)$ values can modify the system’s sensitivity to attacks, potentially enabling mitigation of such attacks. This insight provides a constructive direction for defending SSD. We observed that increasing $Q(k)$ and decreasing $R(k)$ could reduce the system’s sensitivity to attacks. However, this adjustment comes at the cost of a decrease in localization accuracy. While it is possible to improve positioning accuracy by tweaking these parameters, this also makes the system more vulnerable to attacks. Consequently, a dynamic strategy for adjusting $Q(k)$ and $R(k)$ is crucial. This can be achieved through optimization methods or reinforcement learning, which can fine-tune these parameters in real time, balancing between accuracy and security.
8 Conclusion
@@
Acknowledgments
This work is supported by is supported by the National Natural Science Foundation of China (No.62103330, No.62233014), andthe Innovation Foundation for Doctor Dissertation of Northwestern Polytechnical University(CX2023023).
*: Corresponding author

### v2

@@ VI-D Attack Stealthiness
We also use two UAV-specific detectors, NLC [42] and LTW [43], as baselines for comparison. These methods are highly effective at detecting GNSS attacks, and they maintain both strong accuracy and response time, even when identifying targeted stealthy attacks. We follow the threshold in previous work [42]. Fig. 11 demonstrates that in the LTW test group, the detection statistic consistently remains below the threshold boundary, indicating the superior stealth characteristics of SSD attacks. Notably, NLC statistics demonstrate the following three characters (shown in Table III):
@@
These findings demonstrate significant motion-dynamic sensitivity disparities in NLC detectors. However, when attackers combine motion-state transition strategies, positional attacks can realize residual normalization during linear phases and reduce detection sensitivity velocity attacks to counteract statistical fluctuations from nonlinear components. These two methods make SSD show greater stealthiness in hybrid motion missions.
@@
VI-E Attack Effectiveness
We start by determining the optimal combination of $Q_{k}$ and $R_{k}$ for each scenario through offline learning. With these optimal values, we proceed to validate the effectiveness of the SSD. For each combination of INS and trajectory, the UAV is tasked with following the designated path to complete a full mission, while the SSD is deployed to attack the flight. Table IV presents the changes in each metric before and after the attack. On average, ADE/FDE is increased by 425%/591%. The lateral (N)/longitude (E) deviation reaches 3.54/3.46 meters. We will analyze the factor based on the experiment on three scenarios.
@@ VI-E Attack Effectiveness
Normal/Attack (meters) Normal/Attack (meters) Normal/Attack (meters)
ECL EKF2 Trajectory I 20s (0.95)/(5.68) (1.80)/(5.89) (0.89)/(5.66)
CD-EKF Trajectory I (1.62)/(4.64) (1.49)/(7.54) (1.81)/(4.92)
ECL EKF2 Trajectory II 48s (1.31)/(4.09) (1.78)/(6.15) (1.48)/(4.11)
CD-EKF Trajectory II (1.39)/(3.97) (1.99)/(6.47) (2.35)/(5.61)
ECL EKF2 Trajectory III 71s (1.46)/(5.38) (0.95)/(10.47) (1.86)/(6.49)
CD-EKF Trajectory III (1.86)/(6.35) (1.34)/(7.90) (1.66)/(6.73)
@@
(c) Attack visualization in Trajectory III
Figure 12: Attack Visualization
VI-E1 Different Scenarios
@@
In terms of scenarios, the SSD shows a greater increase in positioning error in the purely linear motion state (Fig. 12(a)) compared to the purely nonlinear state (Fig. 12(b)). This difference arises from the time-varying nature of the velocity vector in the nonlinear state, which leads to attenuation of the indirect positional interference caused by the velocity perturbation $G(t_{i},a^{d}_{i};\phi)$. However, experiments with Section VI-D demonstrate that a velocity attack can still ensure the fulfillment of the stealthy precondition. When the mission involves mixed kinematic modes, the combination of velocity and positional attacks results in an additive interaction, as shown in Table IV. The baseline increase is 366.44% for ADE and 1102.11% for FDE, confirming that SSD enhances attack effectiveness synergistically. Next, we will quantitatively analyze the impact of this state on mission completion rates.
VI-E2 Impact on Mission
@@
When the UAV’s trajectory deviation exceeds the tolerance range, the mission completion rate shows a clear decline. In three different scenarios, the APDE metrics increased by $635.96\%$, $277.70\%$, and $348.92\%$, respectively, indicating that SSDs significantly compromise mission reliability. Notably, in Scenario Trajectory III, where the mission duration is 71 seconds, the FDE and APDE deviations reached $10.47m$ and $6.49m$, respectively—the largest among the three scenarios. This demonstrates that SSD lowers the mission completion rate and impedes the UAV’s ability to return accurately after completing the mission (shown in Fig. 12(c)). This effect highlights its kinematic chain impact since the mission-critical point is typically located at the motion state transition (e.g., hover → level flight, acceleration → deceleration). The attack-induced cumulative error in position estimation propagates into the next motion state, where it is further amplified.
VI-E3 Attack Transferability
Different INS architectures manage noise and system uncertainty in distinct ways, potentially impacting the effectiveness of SSD. To assess the cross-system adaptability of the SSD, we conducted validation experiments using CD-EKF and ECL EKF2. CD-EKF relies on continuous-time prediction with discrete-time updates, enhancing its adaptability to errors in nonlinear motion states and making its short-term error estimation more robust. However, SSD effectively exploits CD-EKF’s sensitivity to state uncertainty by inducing motion instability, thereby continuously disrupting navigation accuracy. Experimental results show that under the CD-EKF system, SSD achieves an average improvement of 304.48% in ADE, 473.57% in FDE, and 305.87% in APDE (see Table IV). These findings confirm SSD’s generalizability across EKF-based navigation frameworks.
VII Discussion
@@
This study reveals the coupling mechanism between UAV motion dynamics and cyber attacks’ effectiveness through systematic empirical analysis. Experimental data indicate that changes in motion state can significantly enhance the success rate of attacks. While the SSD approach demonstrates clear advantages, its engineering implementation faces two major challenges: Dependency on A Priori Knowledge. The effectiveness is highly contingent upon the real-time accuracy of the object detection and tracking system. However, the existing YOLOv5 architecture, for instance, exhibits exponential decay in the Intersection over Union (IoU) metric over time in dynamic target tracking scenarios. This results in a tracking failure probability exceeding 73% after 60 seconds of continuous locking. Energy-Concealment Trade-off Paradox. While the sustained attack mode can maintain a stealthiness threshold, it leads to a non-linear increase in energy consumption on the attacking end. This escalation doesn’t align with the requirements in real-world mission scenarios.
@@
In addition, a key assumption of the SSD is that the $Q_{k}$ and $R_{k}$ are preset offline. This is a common practice in current engineering applications. From a theoretical standpoint, adaptive $Q_{k}$ and $R_{k}$ values can modify the system’s sensitivity to attacks, potentially enabling the mitigation of such attacks. This insight provides a constructive direction for defending SSD. We observed that increasing $Q_{k}$ and decreasing $R_{k}$ could reduce the system’s sensitivity to attacks. However, this adjustment comes at the cost of a decrease in localization accuracy. While it is possible to improve positioning accuracy by increasing $R_{k}$ and decreasing $Q_{k}$, this also makes the system more vulnerable to attacks. Consequently, a dynamic strategy for adjusting $Q_{k}$ and $R_{k}$ is crucial. This can be achieved through optimization methods or reinforcement learning, which can fine-tune these parameters in real time, balancing between accuracy and security.
VIII Conclusion
@@
Acknowledgment
This work is supported by the National Natural Science Foundation of China ( No.62233014, No.62103330), and the Innovation Foundation for Doctor Dissertation of Northwestern Polytechnical University (CX2023023).